OpenStudy (anonymous):

Which equation is a quadratic equation? Answer y + 3 = 5x + 10 y + 4x2 = (4x - 2)(x + 3) y - 8x = (x2 + 1)(9x - 7) y = 2x + 7x(x + 1)

OpenStudy (anonymous):

quadratic equations are equations where the highest power of x is 2 , so anything with x^3 in it is not a quadratic. multiply all the brackets out and look for the highest power of x

OpenStudy (anonymous):

is it the 4th one?

OpenStudy (anonymous):

ill go through them one at a time, just a sec

OpenStudy (anonymous):

y + 3 = 5x + 10 well the highest power of x here is 1 , because all we have is an x . this is not a quadratic, and its known as a "linear" equation (because it's a straight line on a graph)

OpenStudy (anonymous):

y + 4x^2 = (4x - 2)(x + 3) y + 4x^2 = 4x^2 + 10x -6 now the 4x^2 terms cancel leaving y = 10x + 6 another linear

OpenStudy (anonymous):

y - 8x = (x2 + 1)(9x - 7) y - 8x = 9x^3 -7x^2 +9x - 7 a cubic!

OpenStudy (anonymous):

finally: y = 2x + 7x(x + 1) y= 2x + 7x^2 +7x the highest power of x is 2 so we have a quadratic like you suggested :D